# aboveskip=2pt, belowskip=2pt

### v1

@@ Appendix B Method Details
B.1 Temporal Sensitivity Analysis
@@ B.2 ArrowRL
Fig. 8 demonstrates ArrowRL’s reward calculation using a VQA example that necessitates causal-temporal reasoning. Here, the fidelity reward ($r_{i}^{\text{fid}}$) alone might be insufficient; for instance, candidates $o_{1}$ and $o_{3}$ both exhibit high similarity to the target $o^{\star}$. The reverse reward distinguishes them by leveraging the semantic difference between forward and reverse video plays. A response $\tilde{o}$ to the reversed video (e.g., describing “organizing or storing the box”) is used as a negative reference. The reverse reward, $r_{i}^{\text{rev}}=1-\text{Similarity}(o_{i},\tilde{o})$, then penalizes forward-video responses like $o_{3}$ if they incorrectly align with the reverse response $\tilde{o}$ (as indicated by red highlighting), thereby favoring temporally aware responses like $o_{1}$ that accurately reflect the forward video’s AoT.
@@ C.1 Quantitative Results
Expanded Ablation Study
Table 6 provides full ablation study results, expanding upon Table 2 in the main paper. Our results validate the design of the reverse reward: its overall contribution (row 2) and dynamic weighting (row 3) are crucial for optimal performance. Note that the reverse reward is not directly applied during training of MCQ-based sequence direction classification; its influence is observed on captioning matching and AoT-VQA tasks (columns 3-5). In addition, the limitations of standard SFT become evident in the direction classification task. As we modify the inference template to be different from training (i.e., by shuffling options), SFT goes through a notable degradation, showcasing its poor generalization under such pertubations, unlike ArrowRL. These ablations collectively demonstrate the importance of ArrowRL’s components for enhancing AoT perception.
Table 6: Ablation study of ArrowRL on AoTBench using Qwen2.5-VL-7B as the base LMM. The effectiveness of ArrowRL’s components is demonstrated: (i) the reverse reward is crucial (removing it in row 2 drops performance below base); (ii) dynamic thresholding for this reward greatly boosts gains (row 3); (iii) SFT fails to generalize on direction classification tasks (row 4).
Model Direc. Cls. Cap. Match AoT- Aver
RFilm UCF T2V V2T VQA age
Qwen2.5-VL-7B [3] 50.0 $\downarrow$ 51.6 $\downarrow$ 53.4 $\downarrow$ 66.6 $\downarrow$ 49.6 $\downarrow$ 56.2 $\downarrow$
+ ArrowRL ($\alpha=0$) 51.4 $\uparrow$ 59.2 $\uparrow$ 49.8 $\downarrow$ 63.7 $\downarrow$ 52.3 $\uparrow$ 55.6 $\downarrow$
+ ArrowRL ($\gamma=1$) 54.9 $\uparrow$ 60.0 $\uparrow$ 50.8 $\downarrow$ 68.4 $\uparrow$ 55.4 $\uparrow$ 58.3 $\uparrow$
+ SFT 49.3 $\downarrow$ 46.2 $\downarrow$ 53.9 $\uparrow$ 66.1 $\downarrow$ 55.3 $\uparrow$ 57.4 $\uparrow$
[…]

### v2

@@ 5 Conclusion
Acknowledgements
This research was supported in part by the UT Austin IFML NSF AI Institute, with compute on the Vista GPU Cluster through the Center for Generative AI (CGAI) and the Texas Advanced Computing Center (TACC). We would like to thank the anonymous reviewers for their thoughtful and constructive feedback, which provided valuable guidance in refining this work.
@@ Appendix B Method Details
B.1 Control Task to Disentangle OOD Effects
To verify that LMMs’ poor performance on AoT tasks (observed in Fig. 1 of the main paper) is not simply due to reversed videos being treated as out-of-distribution (OOD) inputs, we design a control experiment inspired by [3]. Using videos and action labels from UCF101 [58], we formulate a binary-choice question with the prompt, “What action is being performed in this video?” The correct action is provided as one choice, and the incorrect option is randomly drawn from the set of all other action labels. We then evaluate several LMMs on both the original (forward) and time-reversed versions of these videos.
Table 4: We conduct binary action classification on UCF101 as a control task, to disentangle AoT insensitivity from out-of-distribution (OOD) effects. All LMMs achieve near-perfect accuracy regardless of video playback direction, suggesting that failures on AoT tasks are not due to a general inability to process reversed videos.
Model Binary Acc. (%)
forward reverse
GPT-4o [1] 100.0 100.0
Gemini-1.5-Pro [60] 98.2 99.4
LLaVA-OV-7B [34] 99.2 99.4
Qwen2.5-VL-7B [5] 99.4 99.4
The results in Table 4 show that all models perform at near-perfect accuracy on this control task, irrespective of video direction. This contrasts sharply with their near-random performance on our AoT-sensitive sequence direction classification task (Table 1, UCF column). This analysis provides strong evidence that the failure of existing LMMs on AoT tasks is due to their inability to perceive temporal directionality, not a general failure to process reversed videos as OOD inputs.
B.2 Temporal Sensitivity Analysis
@@ B.3 ArrowRL
[…]